Jump to main content.


Drake Chemical

EPA Weekly Situation Reports for December 1998

U. S. ARMY CORPS OF ENGINEERS
DRAKE CHEMICAL SUPERFUND SITE

182 MYRTLE STREET
LOCK HAVEN, PA 17745
PHONE 717-748-4505
FAX 717-748-7549

Wednesday, December 30, 1998

FOR DATES: From 7:00 a.m. Wednesday, December 23, 1998 to 7:00 a.m. Wednesday, December 30, 1998.

The Thermal Destruction Facility (TDF) was on-line for 167.1 hours and processed 4813.8 tons of contaminated soil during the time frame of this report. The TDF was on-line 99.5 % of the time with an average hourly feed rate of 28.8 tons per hour while on-line. The contractor, OHM, continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated approximately 216,197 tons of contaminated soil, which amounts to approximately 76.5 % of the overall contaminated soil on-site.

Construction activities during the time frame of this report included TDF operations, wastewater management, wood shredding, feed preparation, debris separation, lime treatment of soils, ash handling, ash backfilling, dust control to prevent fugitive emissions, dry decontamination and backfilling of rock and concrete debris, and a survey of the excavation bottom. The contractor continued to perform off-site and perimeter air monitoring (PAM) activities.

Excavation continued along the southern end of the exclusion zone, in the general vicinity of the former debris separation building. Excavation activities partially revealed a large nest of discarded drums, drum carcasses and debris. Approximately 80 to 100 drum carcasses were unearthed and separated during the time frame of this report. In addition to the drum carcasses, twenty-three intact drums were also excavated. The initial Health and Safety scan of each drum indicated that the drums contained non-volatile product. The intact drums were wrapped in thick poly (plastic) sheets and stockpiled for overpacking and sampling. The contents of the drums will be sampled and analyzed for chemical reactivity and chemical compatibility. Additional drums and debris are anticipated to be unearthed in this general area as excavation continues toward the north end of the exclusion zone.

Four roll-offs (trailer containers) of non-hazardous PPE and metal debris were shipped off-site to a hazardous waste landfill. One additional roll-off of decontaminated scrap steel was shipped off-site for recycling.

Related site activities included a weekly progress meeting, which was held on Tuesday, December 29, 1998 with representatives from the PADEP, USACE, and OHM in attendance.

Major activities for the next week include: Operation of the TDF, soil feed preparation, excavation, drum handling, lime treatment of soil, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning, and rock backfilling. Nestlerode Contracting will be on-site next week to install a shoring support system for the sheet steel in the exclusion zone.

PREPARED BY: Steve Gibson for Chad Thompson (USACE)


Wednesday, December 23, 1998

FOR DATES: From 7:00 a.m. Wednesday, December 16, 1998 to 7:00 a.m. Wednesday, December 23, 1998.

The Thermal Destruction Facility (TDF) was on-line for 152.7 hours and processed 4381.6 tons of contaminated soil during the time frame of this report. The TDF was on-line 90.9 % of the time with an average hourly feed rate of 28.7 tons per hour while on-line. The contractor, OHM, continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated approximately 212,000 tons of contaminated soil, which amounts to approximately 75 % of the overall contaminated soil on-site. The tonnage and percentage complete is based on the monthly site survey performed during the previous week.

Off-line time to the incinerator during this reporting period was attributed to the stoppage of feed in order to perform additional maintenance to the Dry Ash Drag (DAD) conveyor. Additional maintenance to the DAD was required on Wednesday, December 16th, when a gap was observed to develop between the first two floor plates of the DAD conveyor system. The gap between the plates was allowing ash to drop onto the TDF pad and create minor dust problems in the TDF area. Feed stopped and the DAD was cooled off enough to allow the floor plates to be readjusted. Additional maintenance to the DAD was required on Sunday, December 20th, when feed was stopped in order to remove a piece of a broken idler sprocket from the DAD drive chain and replace the broken idler sprocket and idler shaft. Additional down time was attributed to feed stopping when the master link of the feed screw drive chain broke and had to be replaced on Saturday, December 19th.

The lower hourly average feed rate (less than the target value of 40 tons per hour) is primarily due to the soil material handling capabilities of the dual feed screws when processing wet clayey soil during the winter months. OHM and the USACE are currently investigating alternate soil feed mixes, such as feed mixed with rock, wood or hydrated lime, in order to increase the current feed rate.

OHM's operations crews continued to dismantle the debris separation sprung structure. The skin of the sprung structure was removed and placed in a roll-off for shipment off-site to a hazardous waste landfill. The metal framework of the building was disassembled and decontaminated for future use by OHM.

Construction activities during the time frame of this report included TDF operations and maintenance, wastewater management, wood shredding, feed preparation, debris separation, lime treatment of soils, ash handling, ash backfilling, dust control to prevent fugitive emissions, dry decontamination and backfilling of rock and concrete debris, and a survey of the excavation bottom. The contractor continued electrical work on the lime silo in the exclusion zone. The contractor continued to perform off-site and perimeter air monitoring (PAM) activities.

Excavation continued along the southern end of the exclusion zone, directly south of the former debris separation building location. Excavation activities unearthed several drum carcasses and sections of transite pipe. One large underground storage tank carcass was also unearthed. The tank carcass contained non-reactive site soil. One roll-off of non-hazardous PPE and metal debris was shipped off-site to a hazardous waste landfill.

Related site activities included: * A weekly progress meeting was held on Tuesday, December 22, 1998 with representatives from the USEPA, PADEP, USACE, and OHM in attendance. * The third quarterly stack testing event for metals and particulate emissions was performed on December 17th and 18th 1998. * Kurt Davis (AIR) was on-site to observe stack testing activities from the support zone.

Major activities for the next week include: Operation of the TDF, soil feed preparation, excavation, drum handling, lime treatment of soil, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning, and rock backfilling.

PREPARED BY: Steve Gibson for William Werntges (USACE)


Wednesday, December 16, 1998

FOR DATES: From 7:00 a.m. Wednesday, December 9, 1998 to 7:00 a.m. Wednesday, December 16, 1998.

The Thermal Destruction Facility (TDF) was on-line for 63.3 hours and processed 1678.9 tons of contaminated soil during the time frame of this report. The TDF was on-line 37.7 % of the time with an average hourly feed rate of 26.5 tons per hour while on-line. The contractor, OHM, continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated 207,782.9 tons of contaminated soil, which amounts to approximately 73.5 % of the overall contaminated soil on-site. The tonnage and percentage complete is based on the monthly site survey performed during the previous week.

At the beginning of this reporting period, Wednesday, December 9th, the incinerator remained in a cold shutdown mode in order to perform maintenance to the internal and external components of the TDF. Major maintenance activities included acid washing of the Dry Ash Drag (DAD) cooling plates, replacement of the kiln breech seals and the kiln burners, replacement of the weighbelt conveyance system, replacement of the dual feed screws and the removal of fly ash and slag buildup within the quench tower. The contractor also performed a thorough inspection of the internal refractory of the TDF and performed preventative maintenance to the Wet Ash Drag (WAD) and the DAD conveyors. The monthly interlock checks were performed during the evening hours of Saturday December 12, 1998. Hazardous waste feed was resumed at 0903 hours on Sunday, December 13, 1998.

There was an unplanned Thermal Relief Valve (TRV) opening on Saturday, December 12, 1998 at 0842 hours. The TRV opened when the pressurized cylinder controlling TRV openings decompressed as a result of a pneumatic pipe line becoming blocked with ice. The TRV opening occurred during the incinerator's warming up period, and feed had been off-line for over 6 days. The proper authorities were notified by the contractor and the TRV was closed approximately one hour later.

Off-line time after thermal operations were restarted was primarily attributed to the stoppage of feed in order to perform additional maintenance to the WAD's drive chain and tail shaft. The WAD's drive chain became loose when a sprocket detached from the WAD's tail shaft.

During the cold shutdown of the incinerator, OHM's operations crews continued to implement the close out excavation plan within the exclusion zone. During this period, debris separation equipment was moved to the feed preparation building and the contractor began the disassembling of the debris separation sprung structure. Soil screening activities are now being performed in the feed preparation building.

Construction activities during the time frame of this report included TDF operations and maintenance, wastewater management, feed preparation, debris separation, lime treatment of soils, ash handling, ash backfilling, dust control to prevent fugitive emissions, dry decontamination and backfilling of rock and concrete debris, and a survey of the excavation bottom. The contractor continued electrical work on the lime silo in the exclusion zone. The contractor continued to perform off-site and perimeter air monitoring (PAM) activities.

Excavation continued along the southern end of the exclusion zone, adjacent to the debris separation building. The buildup of sediment in sediment basin 2 was excavated during the time frame of this report. No intact drums were encountered during the time frame of this report and no roll-offs were shipped off-site.

Related site activities included: * A weekly progress meeting was held on Tuesday, December 15, 1998 with representatives from the USEPA, PADEP, USACE, and OHM in attendance. * Kurt Davis (AIR) was given a tour of the Drake Site. * Stack testing for metals and particulate emissions is scheduled for December 17th and 18th, 1998.

Major activities for the next week include: Operation of the TDF, soil feed preparation, excavation, drum handling, lime treatment of soil, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning, and rock backfilling.

PREPARED BY: Steve Gibson for David Modricker (USACE)


Wednesday, December 9, 1998

FOR DATES: From 7:00 a.m. Wednesday, December 2, 1998 to 7:00 a.m. Wednesday, December 9, 1998.

The Thermal Destruction Facility (TDF) was on-line for 115.4 hours and processed 3837 tons of contaminated soil during the time frame of this report. The TDF was on-line 68.7 % of the time with an average hourly feed rate of 33.2 tons per hour while on-line. The contractor, OHM, continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated 205,386.5 tons of contaminated soil, which amounts to approximately 70.1 % of the overall contaminated soil on-site.

Hazardous waste feed to the incinerator was stopped at 0302 hours on Monday, December 7, 1998 in order begin a scheduled cold shutdown maintenance period. During the cold shut down, the contractor performed inspection and preventative maintenance to the dry ash drag conveyor, the wet ash drag conveyor, the weighbelt conveyor and the cyclone. The contractor removed the worn feed screws in preparation for their replacement with a new set of screws. Major maintenance activities for the remainder of the shut down period include the replacement of the kiln burners and the inspection of the refractory material for the whole thermal destruction unit. Hazardous waste feed tentatively scheduled to restart on Sunday, December 13, 1998.

There was a planned Thermal Relief Valve (TRV) opening initiated on Wednesday, December 9, 1998 at 0650 hours. The TRV was opened in order to perform the planned maintenance to the internal components of the incinerator. Feed to the incinerator had been stopped for over 48 hours prior to the opening. The necessary authorities were notified before the TRV opening was initiated.

During the cold shutdown of the incinerator, OHM's operations crews are beginning to implement the close out excavation plan within the exclusion zone. During this period, debris separation equipment will be moved to the feed preparation building and the debris separation sprung structure will be taken down in order to excavate contaminated soils beneath the building's present location. Soil screening activities will now be performed in the feed preparation building.

Construction activities during the time frame of this report included TDF operations and maintenance, wastewater management, feed preparation, debris separation, lime treatment of soils, ash handling, ash backfilling, dust control to prevent fugitive emissions, dry decontamination and backfilling of rock and concrete debris, monthly survey of the exclusion zone and survey of the excavation bottom. The contractor continued electrical work on the lime silo in the exclusion zone. The contractor continued to perform off-site and perimeter air monitoring (PAM) activities.

Excavation continued along the southern end of the exclusion zone, adjacent to the debris separation building. No intact drums were encountered during the time frame of this report. Two roll-offs (trailer containers) of non-hazardous waste was shipped off-site to a hazardous waste landfill. The roll-offs contained PPE, lime bags and miscellaneous metal debris.

Related site activities included: * A weekly progress meeting was held on Tuesday, December 8, 1998 with representatives from the USEPA, PADEP, USACE, and OHM in attendance. * Kurt Davis (AIR) was given a tour of the Drake site in order to observe maintenance activities on the TDF pad. * Stack testing for metals and particulate emissions is scheduled after feed resumes during the third week in December, 1998.

Major activities for the next week include: Operation of the TDF, soil feed preparation, excavation, lime treatment of soil, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning, and rock backfilling.

PREPARED BY: Steve Gibson for David Modricker (USACE)


Wednesday, December 2, 1998

FOR DATES: From 7:00 a.m. Wednesday, November 25, 1998 to 7:00 a.m. Wednesday, December 2, 1998.

The Thermal Destruction Facility (TDF) was on-line for 153.1 hours and processed 5281 tons of contaminated soil during the time frame of this report. The TDF was on-line 91.1 % of the time with an average hourly feed rate of 34.3 tons per hour while on-line. The contractor, OHM, continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated 199706.9 tons of contaminated soil, which amounts to approximately 68.2 % of the overall contaminated soil on-site.

Off-line time during this reporting period is primarily attributed to the stoppage of hazardous waste feed at 0415 hours on Thursday, November 26, 1998 when a conveyor became jammed with a piece of foreign debris. Upon investigation of the incident, the thermal operations crew discovered an obstruction in one of the chains. The contractor removed the metal debris and took additional time to replace a flight in the conveyor. Hazardous waste feed resumed at 1646 hours, once the kiln had reached operating temperature.

Construction activities during the time frame of this report included TDF operations and maintenance, wastewater management, feed preparation, debris separation, lime treatment of soils, ash handling, ash backfilling, dust control to prevent fugitive emissions, dry decontamination and backfilling of rock and concrete debris, surveyed the excavation bottom, and winterization of temperature sensitive components of the TDF. The contractor continued to work on the installation of the lime silo ductwork in the exclusion zone. The contractor continued to perform off-site and perimeter air monitoring (PAM) activities.

Excavation continued along the southeastern end of the exclusion zone adjacent to the debris separation building. A substantial amount of fill with an unusually high moisture content was unearthed in the excavation area. No drums were encountered during the time frame of this report. Two roll-offs (trailer containers) of non-hazardous waste was shipped off-site to a hazardous waste landfill. The roll-off contained PPE, lime bags and miscellaneous metal debris.

Related site activities included: * A weekly progress meeting was held on Tuesday, December 1, 1998 with representatives from the USEPA, PADEP, USACE, and OHM in attendance. * Dr. R. Stehouwer, an Environmental Soil Science Professor from Penn State, met with representatives from the USACE and OHM to discuss the potential and methods for establishing vegetation on the thermally treated soil after remediation is complete at the Drake Site.

Major activities for the next week include: Operation of the TDF, soil feed preparation, excavation, lime treatment of soil, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning, and rock backfilling.

There is an extended maintenance period tentatively scheduled to run from December 7 to December 14, 1998. During this period, the incinerator will be placed in cold shutdown in order to perform maintenance to the internal and external components of the TDF. In addition to the preventative maintenance, the contractor will be performing the monthly calibrations and interlock checks. Hazardous waste feed is scheduled to resume on December 15, 1998, and stack testing is also tentatively scheduled for that week.

PREPARED BY: Ken Hilton (USACE) for David Modricker

Region 3 | Mid-Atlantic Cleanup | Mid-Atlantic Superfund |EPA Home | EPA Superfund Homepage


Local Navigation


Jump to main content.